### v1

# Thermally driven two-sphere microswimmer with internal feedback control

Abstract
We discuss the locomotion of a thermally driven elastic two-sphere microswimmer with internal feedback control that is realized by the position-dependent friction coefficients. In our model, the two spheres are in equilibrium with independent heat baths having different temperatures, causing a heat flow between the two spheres. We generally show that the average velocity of the microswimmer is nonzero when the friction coefficients are position-dependent. Using the method of stochastic thermodynamics, we obtain the entropy production rate and discuss the efficiency of the two-sphere microswimmer. The proposed self-propulsion mechanism highlights the importance of information in active matter and can be a fundamental process in various biological systems.
I Introduction
Microswimmers, such as sperm cells or motile bacteria, are tiny objects moving in viscous environments, and are expected to be relevant to microfluidics and microsystems LaugaBook . By transforming chemical energy into mechanical work, microswimmers change their shapes and move in viscous fluids Hosaka22 . According to Purcell’s scallop theorem, microswimmers in a Newtonian fluid need to undergo non-reciprocal body motion for steady locomotion Purcell1977 ; Ishimoto12 . To realize such a non-reciprocal deformation, various microswimmer models that have more than two degrees of freedom have been proposed Lauga09a . One example is the three-sphere microswimmer in which three in-line spheres are linked by two arms of varying lengths Najafi04 ; Golestanian08 .
Among various generalizations of the three-sphere microswimmer model Yasuda23 , Hosaka et al. considered an elastic three-sphere microswimmer in which the three spheres are in equilibrium with independent heat baths having different temperatures Hosaka17 . It was shown that such a stochastic microswimmer (without any prescribed deformation) can also acquire net locomotion due to thermal fluctuations, and its average velocity is proportional to the heat flow among the spheres Hosaka17 . Later, the average entropy production rate Sou19 ; Sou21 and the time-correlation functions Li24 of the same model were calculated by some of the current authors. In particular, the existence of an antisymmetric part of the cross-correlation function reflects the broken time-reversal symmetry of this microswimmer Li24 .
Prior to the above thermal three-sphere microswimmer model, Kumar et al. proposed a model of an active elastic dimer (AED) in which the friction coefficients of the two spheres depend on their relative coordinate Kumar08 ; Baule08 . Their model is different from the traditional Brownian ratchet models Julicher97 ; Reimann because the motion asymmetry is created internally and is not induced by an external periodic potential. They found that the average velocity of an AED is proportional to the difference between the non-equilibrium noise strengths acting on the two particles Kumar08 ; Baule08 . Such a self-propulsion mechanism can explain, for example, the movement of helicases on DNA Yu06 , the walking of Myosin VI on actin filaments Altman04 , and the collective migration of cell clusters Pages22 .
Given the modern framework of stochastic thermodynamics SekimotoBook ; Jarzynski11 ; Seifert12 and the accumulated knowledge of the thermal three-sphere microswimmer model Hosaka17 ; Sou19 ; Sou21 , the AED model offers a renewed interest, especially when the two spheres have different temperatures without any non-equilibrium noises. Moreover, the internal feedback control considered in the AED model is a crucial mechanism for “informational active matter” that utilizes information (measurement and feedback) instead of energy for various non-equilibrium processes VanSaders24 ; Huang20 . Recently, we have proposed models of Ornstein-Uhlenbeck information swimmers with external and internal feedback controls Hou24 .
In this work, we simplify the original AED model to discuss a thermally driven two-sphere microswimmer with internal feedback control that is realized by the position-dependent friction coefficients. In the current model, the two spheres are in equilibrium with independent heat baths having different temperatures. We show that a combination of heat transfer between the spheres and internal feedback control leads to directional locomotion in the steady state under a noisy condition. We analytically obtain the average velocity and the entropy production rate of this stochastic microswimmer by using stochastic energetics SekimotoBook . With these results, we further estimate the efficiency of the thermal two-sphere microswimmer. We consider that our model explains the fundamental mechanism for transport and locomotion in various active matter and biological systems.
In Sec. II, we explain the model of a thermally driven elastic two-sphere microswimmer with internal feedback control. In Sec. III, we obtain the steady-state probability distribution that is further used to calculate the average velocity and the entropy production rate in Secs. IV and V, respectively. In Sec. VI, we calculate the efficiency of the two-sphere microswimmer by using specific functional forms for the spring potential and the friction coefficients. A brief discussion and summary is given in Sec. VII.
Figure 1: Thermally driven elastic microswimmer consisting of two hard spheres connected by a spring. The positions of the two spheres in a one-dimensional coordinate are denoted by $x_{i}(t)$ ($i=1,2$), and the spring extension is defined by $x=x_{2}-x_{1}$. The spring potential energy is given by $U(x)$, and the two spheres are in equilibrium with independent heat baths having different temperatures $T_{i}$. The friction coefficients $\gamma_{i}(x)$, acting separately on the two spheres, are dependent on the spring extension $x$.
II Model
As schematically shown in Fig. 1, we consider a microswimmer consisting of two hard spheres connected by a spring. The positions of the two spheres in a one-dimensional coordinate are denoted by $x_{i}(t)$ ($i=1,2$), and we further define the spring extension by $x=x_{2}-x_{1}$ that can take both positive and negative values. The spring potential energy is generally given by $U(x)$, which can be a harmonic potential (as we will discuss later) or that of a finitely extensible nonlinear elastic (FENE) spring Baule08 . There are two important assumptions of the model: (i) the two spheres are in equilibrium with independent heat baths having different temperatures $T_{i}$ ($T_{1}\neq T_{2}$) and (ii) the friction coefficients $\gamma_{i}(x)\geq 0$, acting separately on the two spheres, are dependent on the spring extension $x$. To keep the generality of the model as much as possible, we first do not specify the functional forms of $U(x)$ and $\gamma_{i}(x)$.
The coupled overdamped Langevin equations for $x_{i}$ ($i=1,2$) under thermal noise are given by
$\gamma_{i}(x)\dot{x}_{i}(t)=-\partial_{i}U(x)+g_{i}(x)\xi_{i}(t),$ (1)
where $\dot{x}_{i}=dx_{i}/dt$, $\partial_{i}=\partial/\partial x_{i}$ (the summation over $i$ is not taken). The thermal noise $\xi_{i}(t)$ has Gaussian statistics with zero mean and unit variance:
$\langle\xi_{i}(t)\rangle=0,\quad\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime}).$ (2)
According to the fluctuation-dissipation relation, the prefactor of thermal noise is given by $g_{i}(x)=\sqrt{2\gamma_{i}(x)k_{\rm B}T_{i}}$, where $k_{\rm B}$ is the Boltzmann constant DoiBook . The above Langevin equations are multiplicative because the noise is coupled nonlinearly with the stochastic variables Kampenbook . Then, it is necessary to specify the interpretation of the multiplicative noise; otherwise Eq. (1) is meaningless.
Here, we employ the Itô interpretation Gardinerbook , and the Langevin equations in Eq. (1) should be rewritten with the Itô convention $x^{\ast}=x(t)$ as
$\dot{x}_{i}(t)=-\dfrac{\partial_{i}U(x)}{\gamma_{i}(x)}-\dfrac{[g_{i}(x)]^{2}\partial_{i}\gamma_{i}(x)}{2[\gamma_{i}(x)]^{3}}+\dfrac{g_{i}(x^{\ast})}{\gamma_{i}(x^{\ast})}\xi_{i}(t).$ (3)
As pointed out by Lau and Lubensky Lau07 , the second term on the right-hand side is proportional to the temperature, and it is necessary to ensure the proper thermal equilibrium. Such a term reflects thermal fluctuation and generally exists for other interpretations except for the anti-Itô convention Kuroiwa14 . Hereafter, the prime denotes a derivative with respect to the relative coordinate $x=x_{2}-x_{1}$, such as $U^{\prime}(x)=-\partial_{1}U(x)=\partial_{2}U(x)$. Then, the coupled Langevin equations can be explicitly written as
$\dot{x}_{1}(t)$ $=\dfrac{U^{\prime}(x)}{\gamma_{1}(x)}+\dfrac{k_{\mathrm{B}}T_{1}\gamma^{\prime}_{1}(x)}{\gamma^{2}_{1}(x)}+\sqrt{\dfrac{2k_{\mathrm{B}}T_{1}}{\gamma_{1}(x^{\ast})}}\xi_{1}(t),$ (4)
$\dot{x}_{2}(t)$ $=-\dfrac{U^{\prime}(x)}{\gamma_{2}(x)}-\dfrac{k_{\mathrm{B}}T_{2}\gamma^{\prime}_{2}(x)}{\gamma^{2}_{2}(x)}+\sqrt{\dfrac{2k_{\mathrm{B}}T_{2}}{\gamma_{2}(x^{\ast})}}\xi_{2}(t).$ (5)
III Steady-state probability distribution
From Eqs. (4) and (5), the equations of motion for both the center-of-mass coordinate $X=(x_{1}+x_{2})/2$ and the relative coordinate $x=x_{2}-x_{1}$ can be obtained. Since the former equation is a function of the relative coordinate $x$ and the noise only, we first discuss the equation of motion for $x$ in the Itô interpretation given by Baule08 .
$\dot{x}(t)=a(x)+b(x^{\ast})\xi(t).$ (6)
Here, $\xi(t)$ is the superposition of $\xi_{1}(t)$ and $\xi_{2}(t)$ with the same Gaussian white statistics, and the two functions $a(x)$ and $b(x)$ are defined as
$a(x)$ $=-\left(\dfrac{1}{\gamma_{1}(x)}+\dfrac{1}{\gamma_{2}(x)}\right)U^{\prime}(x)+k_{\mathrm{B}}T_{1}\left(\dfrac{1}{\gamma_{1}(x)}+\dfrac{\theta}{\gamma_{2}(x)}\right)^{\prime},$ (7)
$b(x)$ $=\sqrt{2k_{\mathrm{B}}T_{1}\left(\dfrac{1}{\gamma_{1}(x)}+\dfrac{\theta}{\gamma_{2}(x)}\right)},$ (8)
where we have introduced the temperature ratio $\theta=T_{2}/T_{1}$, i.e., $\theta=1$ corresponds to the thermal equilibrium.
The Fokker-Planck equation for the probability distribution $P(x,t)$ associated with the Langevin equation in Eq. (6) can be written in the Itô interpretation as Riskenbook
$\dfrac{\partial}{\partial t}P(x,t)=-\dfrac{\partial}{\partial x}\left[a(x)P(x,t)\right]+\dfrac{1}{2}\dfrac{\partial^{2}}{\partial x^{2}}\left[b^{2}(x)P(x,t)\right].$ (9)
Since we are interested in the steady-state properties of the microswimmer, we obtain the steady-state probability distribution by imposing $\partial_{t}P(x,t)=0$. Moreover, the probability flux can be set to zero since the relative coordinate is bounded by the spring potential $U(x)$. Then, the steady-state probability distribution $P_{\rm s}(x)$ satisfies Baule08
$a(x)P_{\mathrm{s}}(x)-\dfrac{1}{2}\dfrac{\partial}{\partial x}\left[b^{2}(x)P_{\mathrm{s}}(x)\right]=0,$ (10)
and $P_{\rm s}(x)$ can be formally solved as
[…]
where $\mathcal{N}$ is the normalization factor. By substituting Eqs. (7) and (8) into the above expression, we obtain Baule08
[…]
If we further assume that the two friction coefficients are identical for all $x$, i.e., $\gamma_{1}(x)=\gamma_{2}(x)$, the steady-state probability distribution simplifies to
$P_{\mathrm{s}}(x)=\dfrac{\mathcal{N}_{1}}{2k_{\mathrm{B}}T_{1}}\exp\left[-\dfrac{2U(x)}{(1+\theta)k_{\mathrm{B}}T_{1}}\right],$ (13)
which is essentially the Boltzmann distribution modified by the two different temperatures.
IV Average velocity
Having obtained the steady-state distribution function $P_{\mathrm{s}}(x)$ for the relative coordinate $x$, we can calculate the average velocity of the elastic two-sphere microswimmer. Since the center of mass velocity is simply given by $V=\dot{X}=(\dot{x}_{1}+\dot{x}_{2})/2$, we use Eqs. (4) and (5) to obtain the statistical average of $V$ Kumar08 ; Baule08
$\langle V\rangle=\frac{1}{2}\left\langle\left(\dfrac{1}{\gamma_{1}(x)}-\frac{1}{\gamma_{2}(x)}\right)U^{\prime}(x)\right\rangle-\frac{k_{\mathrm{B}}T_{1}}{2}\left\langle\left(\dfrac{1}{\gamma_{1}(x)}-\frac{\theta}{\gamma_{2}(x)}\right)^{\prime}\right\rangle,$ (14)
[…]
When $\gamma_{1}(x)=\gamma_{2}(x)$, the above average velocity further reduces to
$\langle V\rangle=\dfrac{(\theta-1)k_{\mathrm{B}}T_{1}}{2}\left\langle\left(\dfrac{1}{\gamma_{1}(x)}\right)^{\prime}\right\rangle.$ (15)
This result clearly demonstrates that the microswimmer acquires a finite average velocity when $\theta\neq 1$ ($T_{1}\neq T_{2}$) and $\gamma_{1}$ is not constant. In other words, the current microswimmer is driven by thermal energy, and the internal feedback control through $\gamma_{1}(x)$ plays an essential role in the locomotion. However, when $(1/\gamma_{1}(x))^{\prime}$ and $U(x)$ are respectively odd and even functions of $x$, $\langle V\rangle$ vanishes even if the friction coefficient is position-dependent. A similar result to Eqs. (14) and (15) was reported for an AED Kumar08 ; Baule08 , for which it was shown that an asymmetry in the non-equilibrium noise strengths leads to locomotion. In our study, we do not introduce any active fluctuations, which allows us to use the standard framework of stochastic thermodynamics SekimotoBook ; Jarzynski11 ; Seifert12 , as we discuss below.
Figure 2: The average velocity [see Eq. (22)] and the average entropy production rate [see Eq. (23)] of the two-sphere microswimmer when the potential is $U(x)=Kx^{2}/2$ and the friction coefficients are $\gamma_{1}(x)=\gamma_{2}(x)=\gamma[1+\delta\tanh(x/w)]$. Here, we choose $\varepsilon=Kw^{2}/(k_{\mathrm{B}}T_{1})=1$. The filled circles correspond to the result of the numerical simulation. (a) The plot of the dimensionless average velocity $\langle\hat{V}\rangle=\langle V\rangle\gamma/\sqrt{Kk_{\mathrm{B}}T_{1}}$ as a function of $\delta$ for different values of $\theta=T_{2}/T_{1}$. (b) The plot of the dimensionless average entropy production rate $\langle\hat{\dot{\sigma}}\rangle=\langle\dot{\sigma}\rangle\gamma T_{1}/(Kw)^{2}$ as a function of $\delta$ for different values of $\theta$.
Figure 3: The color plot of the efficiency $\eta$ [see Eq. (24)] of the two-sphere microswimmer as a function of $\delta$ and $\theta$. The efficiency $\eta$ vanishes when $\delta=0$.
V Entropy production rate
Next, we calculate the average entropy production rate of the elastic microswimmer in the steady state, which is given by
$\langle\dot{\sigma}\rangle=-\frac{\langle\dot{Q}_{1}\rangle}{T_{1}}-\dfrac{\langle\dot{Q}_{2}\rangle}{T_{2}},$ (16)
where the stochastic heat flow $\dot{Q}_{i}$ is the heat gained by the $i$-th sphere per unit time. According to the stochastic energetics SekimotoBook , it is given by
$\dot{Q}_{i}\,dt=\dfrac{\partial U(x)}{\partial x_{i}}\circ dx_{i},$ (17)
where $\circ$ denotes the Stratonovich product, and the small increment $dx_{i}$ can be obtained from the Langevin equations in Eq. (3). Using the Wong-Zakai theorem to deal with the Stratonovich product and taking the statistical average SekimotoBook , we obtain after some calculation
$\langle\dot{Q}_{i}\rangle$ $=-\left\langle\dfrac{[U^{\prime}(x)]^{2}}{\gamma_{i}(x)}\right\rangle+k_{\mathrm{B}}T_{i}\left\langle\left(\dfrac{U^{\prime}(x)}{\gamma_{i}(x)}\right)^{\prime}\right\rangle.$ (18)
A similar expression was obtained in Eq. (32) of Ref. Cates22 . When the two friction coefficients are constant and identical, the above expression reduces to Eq. (4.27) of Ref. SekimotoBook .
One can easily check that $\langle\dot{Q}_{i}\rangle=0$ when $\theta=1$ by using the equilibrium distribution function in Eq. (12). When $\theta\neq 1$, we consider the total heat flow defined by $\langle\dot{Q}\rangle=\langle\dot{Q}_{1}\rangle+\langle\dot{Q}_{2}\rangle$. By using Eqs. (7) and (8), one can generally show that $\langle\dot{Q}\rangle$ vanishes in the steady state because
$\langle\dot{Q}\rangle=\langle U^{\prime}(x)a(x)\rangle+\dfrac{1}{2}\langle U^{\prime\prime}(x)b^{2}(x)\rangle=0.$ (19)
In the second equality, we have used Eq. (10), which comes from the steady-state condition. Notice that the relation $\langle\dot{Q}\rangle=0$ generally holds even if the two friction coefficients $\gamma_{1}(x)$ and $\gamma_{2}(x)$ are different. Since we have $\langle\dot{Q}_{2}\rangle=-\langle\dot{Q}_{1}\rangle$ from Eq. (19), the average entropy production rate in Eq. (16) can be rewritten as
$\langle\dot{\sigma}\rangle=\frac{1-\theta}{\theta}\frac{\langle\dot{Q}_{1}\rangle}{T_{1}},$ (20)
which vanishes when $\theta=1$. When $T_{1}<T_{2}$ ($\theta>1$), $\langle\dot{Q}_{1}\rangle$ is negative and we have $\langle\dot{\sigma}\rangle>0$, as required from the generalized second law for non-equilibrium systems. A similar statement also holds when $T_{1}>T_{2}$ ($\theta<1$).
VI Example
So far, we have not yet specified the functional forms of the spring potential $U(x)$ and the friction coefficients $\gamma_{i}(x)$. As an explicit example, we choose the harmonic potential energy $U(x)=Kx^{2}/2$, where $K>0$ is the spring constant, and assume that the friction coefficients have the form $\gamma_{1}(x)=\gamma_{2}(x)=\gamma[1+\delta\tanh(x/w)]$, where $\gamma$ is a constant friction coefficient, $\delta$ is a dimensionless parameter satisfying $|\delta|<1$ (recall $\gamma_{i}(x)>0$), and $w$ is a length characterizing the change of the friction coefficients Kumar08 ; Baule08 . With this choice, the steady-state distribution function in Eq. (13) and the average velocity in Eq. (15) become
$P_{\mathrm{s}}(x)=\sqrt{\dfrac{K}{\pi(1+\theta)k_{\mathrm{B}}T_{1}}}\exp\left[-\dfrac{Kx^{2}}{(1+\theta)k_{\mathrm{B}}T_{1}}\right],$ (21)
and
$\langle V\rangle$ $=\dfrac{\delta(1-\theta)}{2\gamma}\sqrt{\dfrac{Kk_{\mathrm{B}}T_{1}}{\pi(1+\theta)}}$
×$\times\int_{-\infty}^{\infty}dz\,\dfrac{\operatorname{sech}^{2}z}{\left(1+\delta\tanh z\right)^{2}}\exp\left(-\dfrac{\varepsilon z^{2}}{1+\theta}\right),$ (22)
respectively, where $\varepsilon=Kw^{2}/(k_{\mathrm{B}}T_{1})$ is a dimensionless parameter. Note that Eq. (21) is simply a Gaussian distribution function. Since the dimensionless integral in Eq. (22) is positive (which will be evaluated numerically below), $\langle V\rangle$ is finite only when $\delta\neq 0$ and $\theta\neq 1$. Moreover, the temperature ratio $\theta$ determines the direction of locomotion, i.e., $\langle V\rangle<0$ when $\delta>0$ and $\theta>1$, or vice versa. Similarly, by calculating the average heat flow $\langle\dot{Q}_{1}\rangle$ in Eq. (18), we obtain the average entropy production rate in Eq. (20) as
$\langle\dot{\sigma}\rangle$ $=\frac{\theta-1}{\theta}\frac{(Kw)^{2}}{\gamma T_{1}}\sqrt{\dfrac{\varepsilon}{\pi(1+\theta)}}\int_{-\infty}^{\infty}dz\,\left[\dfrac{z^{2}}{1+\delta\tanh z}\right.$
$\left.+\dfrac{\delta z\operatorname{sech}^{2}z}{\varepsilon(1+\delta\tanh z)^{2}}-\dfrac{1}{\varepsilon(1+\delta\tanh z)}\right]\exp\left(-\dfrac{\varepsilon z^{2}}{1+\theta}\right),$ (23)
in which the integral is dimensionless.
In Fig. 2(a), we numerically plot the dimensionless average velocity $\langle\hat{V}\rangle=\langle V\rangle\gamma/\sqrt{Kk_{\mathrm{B}}T_{1}}$ in Eq. (22) as a function of $\delta$ for different values of $\theta=T_{2}/T_{1}>1$ when $\varepsilon=1$. To confirm our analytical prediction, we have also performed numerical simulations of the coupled stochastic equations in Eqs. (4) and (5). These Langevin equations were discretized according to the Itô convention. We find a good agreement between the theoretical and numerical results. Notice that $\langle\hat{V}\rangle$ vanishes when either $\theta=1$ or $\delta=0$, as predicted by Eq. (22). Moreover, $\langle\hat{V}\rangle$ is an odd function of $\delta$ because the sign of $\delta$ determines the direction of locomotion. In Fig. 2(b), on the other hand, we numerically plot the dimensionless average entropy production rate $\langle\hat{\dot{\sigma}}\rangle=\langle\dot{\sigma}\rangle\gamma T_{1}/(Kw)^{2}$ in Eq. (23) as a function of $\delta$. For $\theta>1$, $\langle\dot{\sigma}\rangle$ is positive and is an even function of $\delta$. As a result, the entropy production rate is minimized at $\delta=0$ and the corresponding expression is given by $\langle\dot{\sigma}\rangle_{\rm min}=k_{\mathrm{B}}K(\theta-1)^{2}/(2\gamma\theta)\geq 0$.
At this stage, it is worth evaluating the efficiency of the microswimmer defined by Golestanian08 ; Sou21
$\eta=\dfrac{2\gamma\langle V\rangle^{2}}{\langle\dot{\sigma}\rangle T^{\ast}},$ (24)
where $T^{\ast}=(T_{1}+T_{2})/2$ is the average temperature although either $T_{1}$ or $T_{2}$ can also be used. The factor $2\gamma$ in the numerator reflects the fact the microswimmer consists of two spheres with the same average friction coefficient $\gamma$. In Fig. 3, we color-plot the efficiency $\eta$ as a function of $\delta$ and $\theta$. The efficiency $\eta$ vanishes when $\delta=0$ and is an even function of $\delta$ when $\theta$ is fixed. It is worth noting that, in the equilibrium limit of $\theta\rightarrow 1$, the efficiency $\eta$ is finite because both $\langle V\rangle$ and $\langle\dot{\sigma}\rangle$ vanish in this limit [see Eqs. (22) and (23)].
VII Discussion and summary
In our model, the state-dependent friction coefficients play an essential role. The change in the friction coefficient can be caused, for example, by the change in the particle size. In this sense, the current model has a similarity to the model of “pushmepullyou” by Avron et al. Avron05 or the two-sphere model by Pandey et al. Pandey12 . The main difference in our model is that the spheres have different temperatures, and thermal fluctuations cause the locomotion in a stochastic manner. The internal heat flow combined with the feedback control through the position-dependent friction coefficient is the main driving force for maintaining persistent locomotion. In general, there is a contribution to the total entropy production rate coming from the center-of-mass diffusion, which is neglected in Eq. (24Sou21 . In the above efficiency, we have considered the entropy production rate due to the heat flow [see Eq. (16)] that causes the unidirectional motion.
In summary, we have studied the locomotion of a thermally driven elastic two-sphere microswimmer that has position-dependent friction coefficients. Such a microswimmer can acquire nonzero average velocity in the steady state due to the heat flow between the spheres. We have obtained the entropy production rate and further estimated the efficiency of the microswimmer. The proposed self-propulsion mechanism emphasizes the importance of active matter and biological systems driven by information. In the future, we will calculate the time-correlation functions of the microswimmer to quantitatively discuss the degree of broken time-reversal symmetry Li24 .
Acknowledgements.
We thank Z. Xiong for the useful discussion. Z.H.  L.H.  and S.K. acknowledge the support by the National Natural Science Foundation of China (Nos. 12104453, 22273067, 12274098, and 12250710127). Y.H. acknowledges support from JSPS Overseas Research Fellowships (Grant No. 202460086). S.K. acknowledges the startup grant of Wenzhou Institute, University of Chinese Academy of Sciences (No. WIUCASQD2021041). K.Y and S.K. acknowledge the support by the Japan Society for the Promotion of Science (JSPS) Core-to-Core Program “Advanced core-to-core network for the physics of self-organizing active matter” (No. JPJSCCA20230002). J.L. and Z.Z. contributed equally to this work.